OpenStudy (anonymous):

find the measure of an angle if the sum of the measure of its supplement and twice the measure of its compliment is 207 degrees.

OpenStudy (dumbcow):

Let A be the angle. supplement means 180-A, so that A +supplement =180 compliment means 90-A such that A+compliment =90 -> (180-A) + 2(90-A) = 207

OpenStudy (anonymous):

so the angle's measure is?

OpenStudy (dumbcow):

well can you solve a linear equation

OpenStudy (anonymous):

yeah lol.

OpenStudy (dumbcow):

ok then distribute the 2 combine like terms get A by itself

OpenStudy (dumbcow):

try it out, i'll tell you if you are right

OpenStudy (anonymous):

ik i got it angle is 51 degrees

OpenStudy (dumbcow):

yep good job

OpenStudy (anonymous):

thanks :p